OpenStudy (anonymous):

f(x) = 4x + 5 g(x) = 7x - 2 Find: (f + g)(4)

Parth (parthkohli):

\( \color{Black}{\Rightarrow f(4) + g(4)}\) \( \color{Black}{\Rightarrow 4(4) + 5 + 7(4) - 2 }\)

Parth (parthkohli):

Idk what happened to the internet but it ain't showing my LaTeX :|

OpenStudy (anonymous):

f(4) +g(4) = 4*4+5 +7*4 -2 = 47

OpenStudy (anonymous):

f(4)+g(4)=4*4+5+7*4-2=47